OpenStudy (anonymous):

x+1/3=2 can someone show me how this is done?

OpenStudy (anonymous):

is it \(\frac{x+1}{3}=2\) ??

OpenStudy (anonymous):

or \[x+\frac{1}{3}=2\]?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

multiply both sides by 3, then subtract 1 from both sides

OpenStudy (anonymous):

both sides?

OpenStudy (anonymous):

yes, whatever you do to one side, you have to do to the other

OpenStudy (anonymous):

\[\frac{x+1}{3}=2\] \[\frac{x+1}{3}\times 3=2\times 3\] \[x+1=6\]

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

i get it! thank you so much now i'll know how to do the rest of the problems!

OpenStudy (anonymous):

wait 56 isnt the answer right?

OpenStudy (anonymous):

if \(x+1=6\) then \(x=5\)

OpenStudy (anonymous):

6*